OpenStudy (anonymous):

Five added to 2 times a number is equal to 7 less than 4 times the number. What is the number?

OpenStudy (pottersheep):

So let x = the number Five added to 2 times can be expressed as :5 + 2x

OpenStudy (anonymous):

I was thinking it would Be 6?

OpenStudy (pottersheep):

So 5 + 2x = 4x - 7 12 = 2x x =6 Nice you got it!

OpenStudy (uri):

5+2(x)=4(x)-7 ^The first thing you do is convert english to math. @Paradise9541

OpenStudy (anonymous):

Thank you to the both of you! but i go the answer